# Closed-loop Uplink Radio Resource Management in CF-O-RAN Empowered 5G Aerial Corridor

### v1

@@ Abstract
In this paper, we investigate the uplink (UL) radio resource management for 5G aerial corridors with an open-radio access network (O-RAN)-enabled cell-free (CF) massive multiple-input multiple-output (mMIMO) system. Our objective is to maximize the minimum spectral efficiency (SE) by jointly optimizing unmanned aerial vehicle (UAV)–open radio unit (O-RU) association and UL transmit power under quality-of-service (QoS) constraints. Owing to its NP-hard nature, the formulated problem is decomposed into two tractable sub-problems solved via alternating optimization (AO) using two computationally efficient algorithms. We then propose (i) a QoS-driven and multi-connectivity-enabled association algorithm incorporating UAV-centric and O-RU-centric criteria with targeted refinement for weak UAVs, and (ii) a bisection-guided fixed-point power control algorithm achieving global optimality with significantly reduced complexity, hosted as xApp at the near-real-time (near-RT) RAN intelligent controller (RIC) of O-RAN. Solving the resource-allocation problem requires global channel state information (CSI), which incurs substantial measurement and signaling overhead. To mitigate this, we leverage a channel knowledge map (CKM) within the O-RAN non-RT RIC to enable efficient environment-aware CSI inference. Simulation results show that the proposed framework achieves up to 440% improvement in minimum SE, 100% QoS satisfaction and fairness, while reducing runtime by up to 99.7% compared to an interior point solver-based power allocation solution, thereby enabling O-RAN complaint real-time deployment.
@@ I Introduction
To create aerial corridors with guaranteed-coverage , cellular networks must collaboratively provide site-specific beam management and optimized UAV-to-base station associations [18]. The open radio access network (O-RAN) architecture is a compelling enabler, augmenting cellular capabilities with near-real-time control and cross-layer programmability via its RAN intelligent controller (RIC), particularly suited to dynamic aerial UAV [12]. A recent work [4] demonstrates that O-RAN-based closed-loop control jointly optimizing UAV location and transmission directionality delivers approximately 19% uplink (UL) capacity gain while meeting high-definition video quality-of-service (QoS) on multi-cell testbeds. At the RAN-function level, the RIC orchestrates coordinated association and resource allocation, improving energy efficiency and cooperative coverage for scalable aerial corridors [14].
@@
$\bullet$ We propose an O-RAN-enabled CF mMIMO framework integrating a channel knowledge map [19] within the non-real-time (RT) RIC to enable scalable CSI acquisition and circumvent E2 interface limitations.
@@ II System Model
We consider an O-RAN-enabled CF mMIMO system supporting 5G aerial corridor, serving $K$ single-antenna UAVs through $L$ geographically distributed O-RUs, shown in Fig. 1. Each O-RU $\ell\in\mathcal{L}=\{1,2,\ldots,L\}$ is equipped with $N$ antennas and connected to a open distributed/centralized unit (O-DU/O-CU) via fronthaul links. The UAVs operate at altitudes ranging from ground level to several hundred meters and are indexed by $k\in\mathcal{K}=\{1,2,\ldots,K\}$. Let $\mathcal{L}_{k}$ be the set of O-RUs that serve UAV $k$. We assume time-division duplex (TDD) operation, where the channel remains constant over a coherence block of $\tau_{c}$ symbols. Each coherence block is divided into $\tau_{p}$ symbols for UL pilot transmission and $\tau_{c}-\tau_{p}$ symbols for UL data transmission [6].
Figure 1: O-RAN-enabled CF mMIMO system for 5G ariel corridor.
A Walk-through: As illustrated in Fig. 1, the proposed O-RAN–enabled framework comprises two core components: channel knowledge map (CKM) generation [19] and radio resource management (RRM) decision-making. The CKM serves as a repository of propagation statistics indexed by transmitter and receiver locations, thereby enhancing environmental awareness and reducing reliance on real-time CSI acquisition. By mitigating the challenges posed by high-dimensional channels and training overhead, CKM is envisioned as a key enabler for 6G networks demanding extreme capacity, ultra-low latency, and massive connectivity. CKM can be constructed using canonical interpolation methods, Kriging, kernel regression, matrix and tensor completion, deep learning, and environment model–assisted techniques [19]. In the proposed architecture, CKM is implemented as an rApp at the non-RT RIC, while real-time RRM is realized as an xApp at the near-RT RIC. In what follows, we present the operational flow of the proposed framework: (1) The O-RAN network provides feedback and telemetry data to the service management and orchestration (SMO)/O-Cloud via the O1 interface, where it is stored in the global database (DB). (2) The global DB updates the CKM rApp through the R1 interface. (3) The CKM refreshes its information, and the Non-RT RIC supplies policy data (e.g., traffic information, RRM xApp configurations) and updated CSI to the local RAN DB via the A1 interface [^1] RRM algorithms are implemented as xApps at the near-RT RIC. Because the E2 interface supports control signaling but not high-rate CSI exchange, CKM-derived CSI estimates (based on UAV positions obtained via E2 [15]) are delivered through the A1 interface instead.. (4) Using Near-RT internal APIs, the local DB conveys control instructions and CSI to the RRM xApps. (5) The xApps compute RRM decisions and deliver them to the O-DU/O-CU through the E2 interface. (6) Finally, the O-DU/O-CU forwards these decisions to the O-RU via the open fronthaul, which transmits them to UAVs through the air-interface control channels.
@@ II-B Channel Estimation
At the beginning of each coherence block, all $K$ UAVs simultaneously transmit pilot sequences to O-RUs. Like practical system, we consider $K>\tau_{p}$, thereby pilot sequences must be reused among multiple UAVs, leading to pilot contamination. Considering MMSE channel estimation [13] is being used by O-DU/O-CU, the estimate $\widehat{\mathbf{h}}_{k\ell}$ and error $\widetilde{\mathbf{h}}_{k\ell}^{\mathrm{err}}=\mathbf{h}_{k\ell}-\widehat{\mathbf{h}}_{k\ell}$ are independent random variables with $\widehat{\mathbf{h}}_{k\ell}\sim\mathcal{CN}(\overline{\mathbf{h}}_{k\ell},\widehat{\mathbf{C}}_{k\ell})$ and $\widetilde{\mathbf{h}}_{k\ell}^{\mathrm{err}}\sim\mathcal{CN}(\mathbf{0},\mathbf{C}_{k\ell}^{\mathrm{err}})$, where $\widehat{\mathbf{C}}_{k\ell}=\mathbf{C}_{k\ell}-\mathbf{C}_{k\ell}^{\mathrm{err}}$ and $\mathbf{C}_{k\ell}^{\mathrm{err}}=\mathbf{C}_{k\ell}-\mathbf{C}_{k\ell}\bm{\Psi}_{k\ell}^{-1}\mathbf{C}_{k\ell}$ with $\bm{\Psi}_{k\ell}=\tau_{p}^{2}\sum_{i\in\mathcal{P}_{k}}p_{i}^{\mathrm{p}}\mathbf{C}_{i\ell}+\tau_{p}\sigma^{2}\mathbf{I}_{N_{\ell}}$. In $\bm{\Psi}_{k\ell}$, $\mathcal{P}_{k}$ is the set of UAVs sharing UAV $k$’s pilot and $p_{k}^{\mathrm{p}}$ is the pilot power, and $\sigma^{2}$ is the noise power. When $|\mathcal{P}_{k}|>1$, $\widehat{\mathbf{h}}_{k\ell}$ is contaminated by LoS components of interfering UAVs.
@@ III Problem Formulation
Our objective is to jointly optimize the UAV-O-RU association assignment and UL data transmit power allocation for maximizing the minimum SE performance while satisfying scalability and QoS requirements. To this end, we define an association matrix $\bm{A}\in\mathbb{R}^{K\times L}$ where $a_{k\ell}=1$ if UAV $k$ associates with O-RU $l$, and 0 otherwise. Thus, $\mathcal{L}_{k}=\{l:a_{k\ell}=1\}$ with $|\mathcal{L}_{k}|=\sum_{l=1}^{L}a_{k\ell}$. Using these definitions, the SINR in (2) is reformulated as (3). Thus, the joint optimization problem is expressed as follows:
@@
In P0, constraint (4b) ensures every UAV connects to at least one O-RU for ensuring service coverage. Constraint (4c) restricts each O-RU to serve at most $\tau_{p}$ UAVs, dictated by pilot orthogonality, computational capacity, and scheduling overhead [17]. Lastly, constraint (4d) enforces the minimum SE requirement $\text{SE}_{k}^{\min}$ for QoS guarantees. However, P0 is a mixed-integer nonlinear program (MINLP) with binary association $\{a_{k\ell}\}$ and continuous power $\{p_{k}^{\mathrm{u}}\}$ variables. Because the objective (4a) and minimum SE constraint (4d) both involve non-convex $\log_{2}(1+\Gamma_{k})$ functions, where $\Gamma_{k}$ depends nonlinearly on power and association. This renders P0 NP-hard and computationally intractable.
@@ IV-A UAV-O-RU Association Assignment
Since the SINR expression is intricately linked to $\mathbf{A}$ through both channel combining and interference coupling, P1 constitutes a MINLP problem. As finding the global optimum entails prohibitive computational complexity on the order of $\mathcal{O}(2^{KM})$, we develop an efficient heuristic approach to obtain a near-optimal solution, as detailed below.
@@ IV-A1 Proposed UAV-O-RU Association Assignment Scheme
$\bullet$ Stage 1: UAV-centric initialization (Lines 3–7): Each UAV $k$ initially connects to its strongest O-RU $\ell$ based on large-scale fading coefficients, ensuring universal connectivity under the O-RU capacity constraint (4c), thereby satisfies constraint (4b).
@@
     Sort O-RUs by descending $\beta_{k\ell}$: $\mathcal{C}\leftarrow\{\ell_{1},\ldots,\ell_{M}\}\setminus\mathcal{L}_{k}$;
@@ IV-B UL Data Transmit Power Allocation
$\textbf{P1:}\quad\max_{\{p_{k}^{\mathrm{u}}\}\in[0,p^{\max}]}\quad$ $\ \ \min_{k\in\mathcal{K}}\quad\text{SE}_{k},$ (6a)
@@ V Numerical Evaluation
This section evaluates the performance of the proposed UAV–O-RU association and UL transmit power allocation schemes in a O-RAN-enabled CF mMIMO network. Table I summarizes the key simulation parameters following 3GPP UMa-AV specifications [1, 2]. The AO algorithm uses $\epsilon=0.001$ and $I_{\max}=15$ as convergence and iteration limits, respectively. UAV trajectories are assumed predetermined and fixed. Pilot sequences are randomly assigned to UAVs and transmitted at full power. All results are averaged over $500$ Monte Carlo realizations, each with independent O-RU and UAV positions, altitudes, and channel realizations.
@@ V-1 Minimum SE Performance
Fig. 2 demonstrates substantial improvements in the minimum SE performance achieved by the proposed schemes. The joint optimization of association and power allocation (PA + PP or PA + TP) attains up to a 440% gain in minimum SE over the Baseline, highlighting the importance of effectiveness of these two schemes. To isolate individual contributions, we note that the proposed association alone (PA) yields up to an 297.8% improvement over the Baseline, owing to the QoS-driven refinement in Algorithm 1, which dynamically assigns additional serving O-RUs to weak UAVs, a feature that is not included in [17]. As for the proposed power allocation scheme (Algorithm 2), it improves the minimum SE performance by up to 82.6% compared to the Baseline. Notably, the proposed BG-FPPC algorithm (PP) achieves the same minimum SE as the CVX-based solver (TP), confirming its global optimality. This equivalence arises because (a) the outer bisection loop systematically searches for the maximum feasible target SINR $\gamma^{\star}$, and (b) the inner fixed-point iteration precisely computes the minimum power required to meet each target, ensuring that no feasible solution is overlooked. As expected, the minimum SE decreases with increasing UAV density for all schemes due to higher inter-UAV interference. Interestingly, the proposed association schemes (PA + PP and PA + TP) exhibit steeper degradation than the baseline since they proactively assign more O-RUs to QoS-constrained UAVs, thereby increasing interference sensitivity under dense deployments. In contrast, the baseline association remains static and thus less responsive to variations in UAV density.
@@ V-4 Computational Complexity Performance
Fig. 5 illustrates the average runtime for all schemes except the Baseline. The proposed BG-FPPC algorithm (PP) achieves up to 99.1% runtime reduction compared to the CVX-based scheme (TP), with BA + PP outperforming BA + TP by up to 99.9%. Both BA + PP and PA + PP schemes with proposed algorithm (Algorithm 2) achieve 10-150 ms duration, which is within the near-RT RIC time constraint runtime, enabling O-RAN complaint real-time deployment. Notably, PA + PP incurs only 3.4% additional runtime compared to PA + FP, confirming that the AO framework for coordinating the proposed association and power allocation schemes introduces minimal computational overhead.

### v2

@@ Abstract
In this paper, we investigate the uplink (UL) radio resource management for 5G aerial corridors with an open-radio access network (O-RAN)-enabled cell-free (CF) massive multiple-input multiple-output (mMIMO) system. Our objective is to maximize the minimum spectral efficiency (SE) by jointly optimizing unmanned aerial vehicle (UAV)–open radio unit (O-RU) association and UL transmit power under quality-of-service (QoS) constraints. Owing to its NP-hard nature, the formulated problem is decomposed into two tractable sub-problems solved via alternating optimization (AO) using two computationally efficient algorithms. We then propose (i) a QoS-driven and multi-connectivity-enabled association algorithm incorporating UAV-centric and O-RU-centric criteria with targeted refinement for weak UAVs, and (ii) a bisection-guided fixed-point power control algorithm achieving global optimality with significantly reduced complexity, hosted as xApp at the near-real-time (near-RT) RAN intelligent controller (RIC) of O-RAN. Solving the resource-allocation problem requires global channel state information (CSI), which incurs substantial measurement and signaling overhead. To mitigate this, we leverage a channel knowledge map (CKM) within the O-RAN non-RT RIC to enable efficient environment-aware CSI inference. Simulation results show that the proposed framework achieves up to 440% improvement in minimum SE, 100% QoS satisfaction and fairness, while reducing runtime by up to 99.7% compared to an interior point solver-based power allocation solution, thereby enabling O-RAN compliant real-time deployment.
@@ I Introduction
To create aerial corridors with guaranteed coverage , cellular networks must collaboratively provide site-specific beam management and optimized UAV-to-base station associations [18]. The open radio access network (O-RAN) architecture is a compelling enabler, augmenting cellular capabilities with near-real-time control and cross-layer programmability via its RAN intelligent controller (RIC), particularly suited to dynamic UAV scenarios [12]. A recent work [4] demonstrates that O-RAN-based closed-loop control jointly optimizing UAV location and transmission directionality delivers approximately 19% uplink (UL) capacity gain while meeting high-definition video quality-of-service (QoS) on multi-cell testbeds. At the RAN-function level, the RIC orchestrates coordinated association and resource allocation, improving energy efficiency and cooperative coverage for scalable aerial corridors [14].
@@
$\bullet$ We propose an O-RAN-enabled CF mMIMO framework integrating a channel knowledge map [19] within the non-real-time (non-RT) RIC to enable scalable CSI acquisition and circumvent E2 interface limitations.
@@ II System Model
We consider an O-RAN-enabled CF mMIMO system supporting 5G aerial corridor, serving $K$ single-antenna UAVs through $L$ geographically distributed O-RUs, shown in Fig. 1. Each O-RU $\ell\in\mathcal{L}=\{1,2,\ldots,L\}$ is equipped with $N_{\ell}$ antennas and connected to an open distributed/centralized unit (O-DU/O-CU) via fronthaul links. The UAVs operate at altitudes ranging from ground level to several hundred meters and are indexed by $k\in\mathcal{K}=\{1,2,\ldots,K\}$. Let $\mathcal{L}_{k}$ be the set of O-RUs that serve UAV $k$. We assume time-division duplex (TDD) operation, where the channel remains constant over a coherence block of $\tau_{c}$ symbols. Each coherence block is divided into $\tau_{p}$ symbols for UL pilot transmission and $\tau_{c}-\tau_{p}$ symbols for UL data transmission [6].
Figure 1: O-RAN-enabled CF mMIMO system for 5G aerial corridor.
A Walk-through: As illustrated in Fig. 1, the proposed O-RAN-enabled framework comprises two core components: channel knowledge map (CKM) generation [19] and radio resource management (RRM) decision-making. The CKM serves as a repository of propagation statistics indexed by transmitter and receiver locations, thereby enhancing environmental awareness and reducing reliance on real-time CSI acquisition. By mitigating the challenges posed by high-dimensional channels and training overhead, CKM is envisioned as a key enabler for 6G networks demanding extreme capacity, ultra-low latency, and massive connectivity. CKM can be constructed using canonical interpolation methods, Kriging, kernel regression, matrix and tensor completion, deep learning, and environment model–assisted techniques [19]. In the proposed architecture, CKM is implemented as an rApp, a non-RT RIC application, at the non-RT RIC, while real-time RRM is realized as an xApp, a near-RT RIC application, at the near-RT RIC. In what follows, we present the operational flow of the proposed framework: (1) The O-RAN network provides feedback and telemetry data to the service management and orchestration (SMO)/O-Cloud via the O1 interface, where it is stored in the global database (DB). (2) The global DB updates the CKM rApp through the R1 interface. (3) The CKM refreshes its information, and the Non-RT RIC supplies policy data (e.g., traffic information, RRM xApp configurations) and updated CSI to the local RAN DB via the A1 interface [^1] RRM algorithms are implemented as xApps at the near-RT RIC. Because the E2 interface supports control signaling but not high-rate CSI exchange, CKM-derived CSI estimates (based on UAV positions obtained via E2 [15]) are delivered through the A1 interface instead.. (4) Using Near-RT internal APIs, the local DB conveys control instructions and CSI to the RRM xApps. (5) The xApps compute RRM decisions and deliver them to the O-DU/O-CU through the E2 interface. (6) Finally, the O-DU/O-CU forwards these decisions to the O-RU via the open fronthaul, which transmits them to UAVs through the air-interface control channels.
@@ II-B Channel Estimation
At the beginning of each coherence block, all $K$ UAVs simultaneously transmit pilot sequences to O-RUs. As in practical systems, we consider $K>\tau_{p}$, thereby pilot sequences must be reused among multiple UAVs, leading to pilot contamination. Considering MMSE channel estimation [13] is being used by O-DU/O-CU, the estimate $\widehat{\mathbf{h}}_{k\ell}$ and error $\widetilde{\mathbf{h}}_{k\ell}^{\mathrm{err}}=\mathbf{h}_{k\ell}-\widehat{\mathbf{h}}_{k\ell}$ are independent random variables with $\widehat{\mathbf{h}}_{k\ell}\sim\mathcal{CN}(\overline{\mathbf{h}}_{k\ell},\widehat{\mathbf{C}}_{k\ell})$ and $\widetilde{\mathbf{h}}_{k\ell}^{\mathrm{err}}\sim\mathcal{CN}(\mathbf{0},\mathbf{C}_{k\ell}^{\mathrm{err}})$, where $\widehat{\mathbf{C}}_{k\ell}=\mathbf{C}_{k\ell}-\mathbf{C}_{k\ell}^{\mathrm{err}}$ and $\mathbf{C}_{k\ell}^{\mathrm{err}}=\mathbf{C}_{k\ell}-\mathbf{C}_{k\ell}\bm{\Psi}_{k\ell}^{-1}\mathbf{C}_{k\ell}$ with $\bm{\Psi}_{k\ell}=\tau_{p}^{2}\sum_{i\in\mathcal{P}_{k}}p_{i}^{\mathrm{p}}\mathbf{C}_{i\ell}+\tau_{p}\sigma^{2}\mathbf{I}_{N_{\ell}}$. In $\bm{\Psi}_{k\ell}$, $\mathcal{P}_{k}$ is the set of UAVs sharing UAV $k$’s pilot and $p_{k}^{\mathrm{p}}$ is the pilot power, and $\sigma^{2}$ is the noise power. When $|\mathcal{P}_{k}|>1$, $\widehat{\mathbf{h}}_{k\ell}$ is contaminated by LoS components of interfering UAVs.
@@ III Problem Formulation
Our objective is to jointly optimize the UAV-O-RU association assignment and UL data transmit power allocation for maximizing the minimum SE performance while satisfying scalability and QoS requirements. To this end, we define an association matrix $\bm{A}\in\mathbb{R}^{K\times L}$ where $a_{k\ell}=1$ if UAV $k$ associates with O-RU $\ell$, and 0 otherwise. Thus, $\mathcal{L}_{k}=\{\ell:a_{k\ell}=1\}$ with $|\mathcal{L}_{k}|=\sum_{\ell=1}^{L}a_{k\ell}$. Using these definitions, the SINR in (2) is reformulated as (3). Thus, the joint optimization problem is expressed as follows:
@@
In P0, constraint (4b) ensures every UAV connects to at least one O-RU for ensuring service coverage. Constraint (4c) restricts each O-RU to serve at most $\tau_{p}$ UAVs, dictated by pilot orthogonality, computational capacity, and scheduling overhead [17]. Lastly, constraint (4d) enforces the minimum SE requirement $\text{SE}_{k}^{\min}$ for QoS guarantees. However, P0 is a mixed-integer nonlinear programming (MINLP) with binary association $\{a_{k\ell}\}$ and continuous power $\{p_{k}^{\mathrm{u}}\}$ variables. Because the objective (4a) and minimum SE constraint (4d) both involve non-convex $\log_{2}(1+\Gamma_{k})$ functions, where $\Gamma_{k}$ depends nonlinearly on power and association. This renders P0 NP-hard and computationally intractable.
@@ IV-A UAV-O-RU Association Assignment
Since the SINR expression is intricately linked to $\mathbf{A}$ through both channel combining and interference coupling, P1 constitutes a MINLP. As finding the global optimum entails prohibitive computational complexity on the order of $\mathcal{O}(2^{KL})$, we develop an efficient heuristic approach to obtain a near-optimal solution, as detailed below.
@@ IV-A1 Proposed UAV-O-RU Association Assignment Scheme
$\bullet$ Stage 1: UAV-centric initialization (Lines 3–7): Each UAV $k$ initially connects to its strongest O-RU $\ell$ based on large-scale fading coefficients, ensuring universal connectivity under the O-RU capacity constraint (4c), thereby satisfying constraint (4b).
@@
     Sort O-RUs by descending $\beta_{k\ell}$: $\mathcal{C}\leftarrow\{\ell_{1},\ldots,\ell_{L}\}\setminus\mathcal{L}_{k}$;
@@ IV-B UL Data Transmit Power Allocation
$\textbf{P2:}\quad\max_{\{p_{k}^{\mathrm{u}}\}\in[0,p^{\max}]}\quad$ $\ \ \min_{k\in\mathcal{K}}\quad\text{SE}_{k},$ (6a)
@@ V Numerical Evaluation
This section evaluates the performance of the proposed UAV–O-RU association and UL transmit power allocation schemes in an O-RAN-enabled CF mMIMO network. Table I summarizes the key simulation parameters following 3GPP UMa-AV specifications [1, 2]. The AO algorithm uses $\epsilon=0.001$ and $I_{\max}=15$ as convergence and iteration limits, respectively. UAV trajectories are assumed predetermined and fixed. Pilot sequences are randomly assigned to UAVs and transmitted at full power. All results are averaged over $500$ Monte Carlo realizations, each with independent O-RU and UAV positions, altitudes, and channel realizations.
@@ V-1 Minimum SE Performance
Fig. 2 demonstrates substantial improvements in the minimum SE performance achieved by the proposed schemes. The joint optimization of association and power allocation (PA + PP or PA + TP) attains up to a 440% gain in minimum SE over the Baseline, highlighting the effectiveness of these two schemes. To isolate individual contributions, we note that the proposed association alone (PA) yields up to an 297.8% improvement over the Baseline, owing to the QoS-driven refinement in Algorithm 1, which dynamically assigns additional serving O-RUs to weak UAVs, a feature that is not included in [17]. As for the proposed power allocation scheme (Algorithm 2), it improves the minimum SE performance by up to 82.6% compared to the Baseline. Notably, the proposed BG-FPPC algorithm (PP) achieves the same minimum SE as the CVX-based solver (TP), confirming its global optimality. This equivalence arises because (a) the outer bisection loop systematically searches for the maximum feasible target SINR $\gamma^{\star}$, and (b) the inner fixed-point iteration precisely computes the minimum power required to meet each target, ensuring that no feasible solution is overlooked. As expected, the minimum SE decreases with increasing UAV density for all schemes due to higher inter-UAV interference. Interestingly, the proposed association schemes (PA + PP and PA + TP) exhibit steeper degradation than the baseline since they proactively assign more O-RUs to QoS-constrained UAVs, thereby increasing interference sensitivity under dense deployments. In contrast, the baseline association remains static and thus less responsive to variations in UAV density.
@@ V-4 Computational Complexity Performance
Fig. 5 illustrates the average runtime for all schemes except the Baseline. The proposed BG-FPPC algorithm (PP) achieves up to 99.1% runtime reduction compared to the CVX-based scheme (TP), with BA + PP outperforming BA + TP by up to 99.9%. Both BA + PP and PA + PP schemes with proposed algorithm (Algorithm 2) achieve 10-150 ms duration, which is within the near-RT RIC time constraint runtime, enabling O-RAN compliant real-time deployment. Notably, PA + PP incurs only 3.4% additional runtime compared to PA + FP, confirming that the AO framework for coordinating the proposed association and power allocation schemes introduces minimal computational overhead.